OpenStudy (anonymous):

Simplify completely: (5x^2-21x-20)/(5x^2-16x-16)

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

How much longer guys?

OpenStudy (cwrw238):

factor top and bottom and see if u can cancel 5x^2-21x-20 = (5x )(x ) can you complete this? you need to numbers whose product is -20 and when one is multiplied by 5 and the other added will give -21

OpenStudy (anonymous):

I just need the bottom factorf please.

OpenStudy (anonymous):

I know how, it just take me forever.

OpenStudy (cwrw238):

5x^2-16x-16 = (5x + 4)(x - ) can u fill in the blank

OpenStudy (cwrw238):

5 times some value + 4 = -16 - thats not too hard

OpenStudy (cwrw238):

5x + 4 = -16 whats x?

OpenStudy (anonymous):

-4=x